# A Comprehensive Survey of Dataset Distillation

### v2

@@ 8.2 Future Directions
Although the existing DD methods have shown impressive results in compressing massive datasets, there are still many areas that are worth exploring to promote dataset distillation. In this section, we will discuss some promising directions that shed light on future studies.
Fine-tune with synthetic data. With the rapid development of foundational models [7], the mainstream deep learning pipeline gradually converts from scratch learning to pretrained learning or transfer learning. The foundational models of BERT [27], GPT [92, 93, 8], etc., are trained with billions of data points to learn an appropriate latent representation. With only fine-tuning on specific datasets, the foundational models can achieve extraordinary performance on various downstream tasks and outperform models learned from scratch by a large margin. To this end, conducting dataset distillation in the fine-tuning regime is a promising direction to cooperate with foundational models. With synthetic data, the fine-tuning process will be faster while the downstream performance is preserved. Despite possessing a similar goal, distilling datasets based on pretrained models might be more challenging than the initialized models because the pretrained models are harder to collect for repeated distillation to enhance synthetic datasets’ robustness w.r.t. model initialization. Apart from condensing datasets in the fine-tuning stage, it is also an interesting topic to distil datasets in the vanilla training stage so that models trained on synthetic datasets can be easier or faster to fine-tune by specific datasets.
Initialization augmentation. Benefiting from excellent efficiency and compatibility, data augmentation has become an essential strategy for enhancing the model’s generalization w.r.t. data distribution [102, 15, 88]. From basic cropping and flipping [121] to the mixup method [123], data augmentation efficiently constructs variants of training images with prior rules to improve the model’s robustness to patterns. In dataset distillation, synthetic datasets are also sensitive to the parameter initialization and network architecture adopted in the distillation process [114, 129]. Existing works alleviate this problem by trivial repeated distillation w.r.t. different initializations and then taking an expectation [132]. However, the random selection of initialization is not an elaborate strategy and has limited improvement. Zhang et al. [124] showed that employing early-stage models trained with a few epochs as the initialization can achieve better distillation performance, and they further proposed weight perturbation methods to efficiently generate early-stage models for repeated distillation. Therefore, it is important for many dataset distillation algorithms to investigate how to design the initialization augmentation so that synthetic datasets achieve better generalization to different initializations and architectures.
Preprocessing of target datasets. To better extract knowledge from training data, preprocessing, such as normalization, is a common approach that reshapes the structure of data by well-designed mappings [32]. In existing dataset distillation methods, only a few works discuss the influence of preprocessing on final distillation performance [86]. Nguyen et al. [86, 87] and Cazenavette et al. [13] preprocessed target datasets with zero-phase compenent analysis (ZCA) whitening [53] before distillation and achieved satisfying empirical results. Through elaborate preprocessing, more explicit information can emerge, and the target datasets might be easier to distill. In addition, this preprocessing is compatible with existing dataset distillation algorithms and thus can be deployed without extra effort. For this reason, it is useful to investigate the preprocessing of target datasets for distillation acceleration and also performance improvement.

### v3

@@ 9.2 Future Directions
Despite aforementioned challenges existed in DD, there are still many promising directions that are worth exploring to (1) enhance dataset distillation performance and (2) develop trustworthy dataset distillation.
Optimization objective of DD. As discussed in Sections 5 and 6.1, optimizing factorized codes and decoders that cooperatively generate synthetic data outperform the direct optimization by a large margin. This suggests that the selection of optimization objectives (e.g., RGB pixels, latent codes and decoders) has a significant influence on the final distillation performance. Apart from optimization on original data or factorized spaces, Wang et al. [99] turned to optimize the generative model to synthesize unlimited informative data, and Cazenavette et al. [53] optimized the intermediate features of generative models to generate more realistic synthetic data for better cross-architecture generalization. Therefore, investigating optimization on proper information carriers, such as optimizing a few 3D data to summarize information of 2D images, is a promising direction for further DD performance improvement.
Model augmentation. During the process of DD, neural networks within different training stages are involved to compute the distillation loss, and then the synthetic data are accordingly updated. This triggers a problem that distilled data are sensitive to network parameters, and causes unsatisfied DD performance when training new networks whose parameters have a large discrepancy to models used for calculating distillation loss. Therefore, the sensitivity to model parameters is closely related to DD performance. To alleviate this sensitivity, early works use a simple model augmentation strategy that conducts DD on networks with different initialization and training stages [11, 24]. Recently, Zhang et al. [100] showed that employing early-stage models trained with a few epochs as the initialization can achieve better distillation performance, and they further proposed weight perturbation methods to efficiently generate early-stage models for calculating distillation loss. Therefore, it is important for dataset distillation to investigate model augmentation algorithms to reduce the sensitivity of distilled data to model parameters and further improve DD performance.
Cross-architecture transferability. Because the distillation of synthetic data is closely related to the specific network architecture (or kernels in the KRR approach), there is naturally a performance drop when distilled datasets are applied to train networks with other unseen architectures. Recently, Cazenavette et al. [53] argued that this low cross-architecture transferability comes from direct optimization on the pixel space, which makes the distilled data overfit to specific architectures used in the DD process. With this argument, they regularized the distillation by optimizing the latent codes that are fed into pretrained generative models for distilled data generation, which largely increases the cross-architecture performance due to more realistic synthetic data. A decent cross-architecture generalization is significant and allows synthetic data to train more complex models that are prohibitively used in DD process due to unbearable computing overhead.
Private dataset distillation. As the real data are involved in dataset distillation, it is necessary to analyze the distilled data from the perspective of privacy protection.
Theoretically, Dong et al. [101] and Zheng and Li [102] built the connection between dataset distillation and differential privacy (DP), and proved the superiority of dataset distillation in privacy preservation over conventional private data generation methods. However, Carlini et al. [103] argued that there exist flaws in both experimental and theoretical evidence in [101]: (1) membership inference baseline should include all training examples, while only $1\%$ of training points are used to match the synthetic sample size, which thus contributes to the high attack accuracy of baseline; and (2) the privacy analysis is based on an assumption stating that the output of learning algorithms follows a exponential distribution in terms of loss, while the assumption has already implied differential privacy. Empirically, Chen et al. [104] employed gradient matching to distill private datasets by adding DP noise to the gradients of synthetic data, which was followed by Vinaroz and Park [105] who privatized KIP with DP-SGD to achieve better privacy-accuracy trade-off compared to [104]. Therefore, a powerful private dataset distillation algorithm helps reassuringly publish distilled data and alleviate data abuse.
Robust dataset distillation. It is significant to develop robust DD algorithms such that models trained on distilled data own satisfied robustness w.r.t. malicious attacks and ambiguous test data, or on the contrary, attack downstream models via manipulating the synthetic data.
To enhance adversarial robustness, Tsilivis et al. [106] and Wu et al. [107] employed dataset distillation to extract the information of adversarial examples and generate robust datasets. Then standard network training on the distilled robust dataset is sufficient to achieve satisfactory robustness to adversarial perturbations, which substantially saves computing resources compared to expensive adversarial training. For backdoor attack, Liu et al. [108] considered injecting backdoor triggers into the small distilled dataset. Because the synthetic data has a small size, direct trigger insertion is less effective and also perceptible. Therefore, these coworkers turned to insert triggers into the target dataset during the dataset distillation procedure. In addition, they propose to iteratively optimize the triggers during the distillation process to preserve the triggers’ information for better backdoor attack performance. These defense and attack algorithms in terms of DD invoke the distillation of robust synthetic datasets, and models trained on them can better defend against various attacks.
Recently, Zhu et al. [109] revealed that models trained on distilled data tend to output over-confident predictions and have a bad calibration, due to discarding semantically meaningful information. To tackle this problem, they propose to randomly mask the synthetic data to compel them to contain more semantically complete information during the distillation process. Apart from estimating proper uncertainty for in-distribution data, it is also important to detect out-of-distribution (OoD) examples for models deployed in the open world. Ma et al. [110] additionally synthesized outliers based on corruption transformations of real data, which are called pseudo-outliers. Then the synthetic outliers are encompassed to train downstream models for enhancing the capability of OoD detection. Therefore, it is significant for trustworthy learning to construct distilled datasets that can induce proper uncertainty w.r.t. both in-distribution and out-of-distribution test data.